Help four problems solving equations with variables on each side with fractions. Thank you
4/5y - 8 = 2/5y +16
x/0.4=2x +1.2
3/4k + 16 = 2 - 1/8k
1/3b + 8 = 1/2b -4

Answer by stanbon(75887)

(4/5)y - 8 = (2/5)y +16
----
(2/5)y = 24
(1/5)y = 12
y = 60
---------------------------
x/0.4 = 2x +1.2
Multiply thru by 0.4 to get:
x = 0.8x + 0.48
0.2x = 0.48
20x = 48
x = 2.4
---------
(3/4)k + 16 = 2 - (1/8)k
Multiply thru by 8 to get:
6k + 8*16 = 16 - k
7k = -7*16
k = -16
==============
Cheers,
Stan H.
=================
1/3b + 8 = 1/2b -4